OpenStudy (anonymous):

Graph the equation using the slope and y-intercept. x+y=0

OpenStudy (anonymous):

try to consider y= mx+b

OpenStudy (anonymous):

Im not sure if the slope would be 1 and the y-intercept would be 0? :/

OpenStudy (anonymous):

y int is correct, but the slope isn't. try again

OpenStudy (anonymous):

I dont know then. My teacher said if there was just x it had an imaginary 1 in front of it. :P

OpenStudy (anonymous):

yes that's true, but your equation isn't in terms of y = what do you have to do to x?

OpenStudy (anonymous):

to solve for y-intercept substitute the value of x to 0 and solve for y eg// x+y=0 0+y=0 y=0 y- intercept is 0

OpenStudy (anonymous):

Ok I have the y-intercept down, but Im confused with how to get the slope

OpenStudy (anonymous):

x+y=0 y= -x+ 0 from the slope-intercept equation y=mx+b m stands for slope

OpenStudy (anonymous):

so your slope would be?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

yes, right. Go graph it now :D

OpenStudy (anonymous):

Okay thanks :D